What is the mass of
10 moles of sodium sulphite (Na2SO3)?

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

10 moles of sodium sulphite (Na2SO3)
Mass of 1 mole of Na2SO3= 2 × 23 + 32 + 3 ×16
= 46 + 32 + 48 = 126 g
Mass of 10 moles of Na2SO3= 126 × 10
= 1260 g

Thanks